If 7 is subtracted from twice a number, the result is 5 increased by 4. Find the number.
Answer by richwmiller(17219)

let x=the number
2x-7=5+4
2x-7=9
add 7 to both sides
2x=16
x=8
check
16-7=9
16=16
ok